# Analytic Models for the Capacity Distribution in MDG-impaired Optical SDM Transmission

### v2

@@ I Introduction
The increased capacity of multi-mode or multi-core fibers is a field of active research to handle the growing data traffic [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, unequal gains in amplifiers or unequal losses in passive devices and fibers generate an effect known as mode-dependent gain (MDG) and mode-dependent loss (MDL). In systems with MDG or MDL [11], the spatial coupling during fiber propagation turns the channel capacity into a random variable, introducing a probability of outage [12]. Communications systems with a stochastic signal-to-noise ratio (SNR) need to be designed to maintain the outage probability below accepted levels. To properly assess outages in the design of future space-division multiplexing (SDM) systems, statistical models for the capacity are necessary.
@@
The capacity loss statistics for a set of amplification schemes are discussed in [13]. Expressions for the variance of the per-mode capacity loss for four different amplification schemes are derived. However, both per-mode capacity and total capacity distributions are not analyzed. Distributions for the per-mode modal gains have been presented in [14]. In addition, the paper obtains a closed-form expression for the average capacity. However, the capacity variance is not calculated. The impact of frequency diversity on the capacity distribution has been discussed in [15], but no analytical expression for the capacity distribution has been presented. The exact distribution for a single-mode-fiber optical system with polarization-dependent gain (PDG, $D=2$) has been presented in [16]. However, the statistics of SDM systems with higher mode counts have been addressed by simulation.
@@
In this paper, we provide analytical expressions for the variance of the capacity distribution [^1] An analytical solution for the total capacity mean is presented in [14, Eq. (10)].. The derivations are based on the assumption of a Gaussian distribution for the per-mode and total capacity probability density functions (PDFs). The Gaussian assumption for the total capacity distribution has been discussed in [17], confirming its validity for increased mode counts. The joint capacity distribution is approximated as a multivariate normal, and the total capacity is obtained as the linear combination of the individuals. The derivations rely on the modal gain distributions outlined in [14].
II Instantaneous Capacity
Figure 1: (a) Logarithmic gain $\lambda_{\mathrm{dB}}$ distribution of each mode and ensemble, considering the sum of the linear gains equals to $D$. The ensemble distribution is normalized by $D$ in order to facilitate visual analysis. The per-mode distributions are obtained via simulation for ${D=6}$ and ${\sigma_{mdg}=5~{}\mathrm{dB}}$. (b) Instantaneous capacity distribution of each mode and total. The distributions are obtained via simulation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
@@ II Instantaneous Capacity
where SNR is the ratio of total signal power and total noise power, $\mathbf{H}$ is the channel transfer matrix, and $\mathbf{I}$ is the identity matrix of size $D$, where $D$ is the mode count. The total instantaneous capacity can also be obtained from the eigenvalues $\lambda$ of the $\mathbf{H}\mathbf{H}^{H}$ [18] operator
@@
where $\lambda_{i}$ is the $i^{th}$ mode power gain experienced in the system. We consider noise as spatially white [^2] The spatial whiteness assumption has stronger validity for higher mode counts and wider bandwidths [19]., yielding equal noise power for all modes [14]. Assuming that SDM amplifiers maintain a constant total power after amplification, the sum of all $\lambda_{i}$ is set to $D$.
Equation (2) provides the instantaneous capacity for a single channel realization. In MDG-impaired systems, the random modal drift during fiber propagation results in a time-varying matrix $\mathbf{H}$, which in turn leads to fluctuations in both per-mode and total capacities. As an example, the modal-gain distribution for an optical transmission link affected by MDG is illustrated in Fig. 1(a). These gains result in a time-varying per-mode and total capacities that are illustrated in in Fig. 1(b). In [14], Ho et al. resort to random matrix theory to demonstrate that the overall PDF of an ensemble of all $\lambda_{i}$ eigenvalues, in decibels, is given by the well-known Gaussian unitary ensemble (GUE) spectral distribution, shown in Fig. 1(a) for $D=6$ [20, 21]. However, obtaining closed-form analytical models for the individual marginal $\lambda_{i}$ eigenvalue distributions is not trivial, except for specific cases, such as low eigenvalue counts (${D\leq 3}$[22], or for the smallest and largest eigenvalues as specified by the Tracy-Widom distribution [23].
III Per-Mode Individual Capacity Distributions
@@
We evaluate the statistics of the total capacity $C_{T}$ based on the per-mode individual $\lambda_{i}$ and $C_{i}$ distributions. Approximations on $C_{i}$ distributions are obtained from two methods derived from the GUE spectral and Wigner semicircular distribution for $\lambda_{i}$.
@@ III-A Approximations on the $C_{i}$ distribution based on the GUE spectral distribution for $\lambda_{i}$
The distribution of the $\lambda_{\mathrm{dB}}$ eigenvalues for a $D^{th}$ order GUE is given by [14, Table 1], [22, Eq. (S.15)]
@@
where $\alpha_{\lambda_{\mathrm{dB}},D}$ is a normalization factor such that the PDF integral equals 1, and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are polynomial coefficients. $\alpha_{\lambda_{\mathrm{dB}},D}$ and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are constant for a given number of modes. For $D\leq 8$ the values of $\alpha_{\lambda_{\mathrm{dB}},D}$ and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are provided in [14, Table 1], and the derivation of the coefficients is presented in appendix A. Values [^3] The notation $\sigma_{\mathrm{mdg}}$ was preferred to $\sigma_{\lambda_{\mathrm{dB}}}$, as the former is predominant in literature. $\sigma_{\mathrm{mdg}}$ and $\mu_{\lambda_{\mathrm{dB}}}$ are the standard deviation and the mean of $f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})$, respectively. Value $\mu_{\lambda_{\mathrm{dB}}}$ is given such that the linearised eigenvalue distribution $f_{\lambda}(\lambda)$ has unitary mean, and is given by (see appendix B)
@@
Figure 2: Analytical approximation of per-mode capacity mean (a) and standard deviation (b) for ${\mathrm{SNR}=10~{}\mathrm{dB}}$ and ${\sigma_{mdg}=5~{}\mathrm{dB}}$. For ${D\leq 8}$ the GUE distribution method is considered. Higher mode counts use the Wigner semicircular CDF method.
Figure 3: Gaussian per-mode capacity PDF s for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
@@
An evaluation of $\mu_{C_{i}}$ and $\sigma_{C_{i}}$ for $2\leq D\leq 8$ is presented in Fig. 2, exhibiting an excellent agreement. The accuracy of the Gaussian approximation for the individual capacity distributions $f_{C_{i}}$, for a system with $D=6$, $\mathrm{SNR}=10~{}\mathrm{dB}$ and $\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}$, is shown in Fig. 3. The results reveal, again, a suitable agreement between simulation and the derived analytical curves.
@@
The limiting distribution of $\lambda_{\mathrm{dB}}$ as $D\rightarrow\infty$ is given by the Wigner semicircular distribution as [20, I][14, Table 1]
@@
An assessment of the CDF-derived approximation can be found in Fig. 4, for a system with $D=6$, $\mathrm{SNR}=10~{}\mathrm{dB}$ and $\sigma_{mdg}=5~{}\mathrm{dB}$, showing excellent agreement.
@@
IV Total Capacity Distribution
@@
Alternatively, an already proposed exact solution for $\mu_{C_{T}}$ has been proposed in [14, Eq. (10)][16, Eq. (12)]
@@
Figure 5: Correlation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.
@@
$\rho_{i,j}=e^{-|i-j|}+\left(e^{-|i-j|}-1\right)\left(\gamma_{0}+\gamma_{1}\cdot\sigma_{mdg}^{2.75}\right),$ (21)
@@
$\left(\gamma_{0},\gamma_{1}\right)=\mathrm{argmin}_{\gamma_{0},\gamma_{1}}\left(\mathbb{E}\left\{\left[\mathrm{ln}\left(\widehat{\sigma}^{2}_{C_{T}}\right)-\mathrm{ln}\left(\sigma^{2}_{C_{T}}\right)\right]^{2}\right\}\right).$ (22)
The methodology to obtain the coefficients is performed iteratively until a tolerance threshold is reached. The algorithm implemented in this work firstly updates $\gamma_{0}$ until the error for the lowest considered $\sigma_{\mathrm{mdg}}$ is smaller than a defined tolerance. The value of $\gamma_{1}$ is updated considering the MSLE for the entire evaluated $\sigma_{\mathrm{mdg}}$ range, with small corrections to $\gamma_{0}$ if the resulting approximation function is not monotonic increasing. We select the logarithmic error because the capacity variance scales exponentially with $\sigma_{mdg}$ when measured in decibels. The coefficients for a practical range of mode counts and SNR are given in Table I.
An evaluation of the analytical method is presented in Fig. 6. The results indicate an accurate variance approximation for SNR values lower than $20~{}\mathrm{dB}$ for all evaluated mode counts. For $\textrm{SNR}=20~{}\mathrm{dB}$, there is a higher divergence, particularly with the Wigner semicircular CDF method. The source of estimation discrepancy likely stems from the per-mode individual distributions estimated from the capacity CDF, as well as approximations made in the correlation function. Fig. 7 evaluates the accuracy of the method based on Wigner semicircular CDF for different values of $\sigma_{mdg}$ and mode counts. The results evidence accurate estimates for mode counts higher than 20. Fig. 8 compares the simulated total capacity distribution with the analytical distributions produced with the Wigner semicircular CDF method. The presented Gaussian approximation exhibits a good agreement with the simulated curve.
Figure 8: Total Gaussian PDF for ${D=20}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$.

### v3

@@ I Introduction
The increased capacity of multi-mode or multi-core fibers is a field of active research to handle the growing data traffic [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, unequal gains in amplifiers or unequal losses in passive devices and fibers generate an effect known as mode-dependent gain (MDG) and mode-dependent loss (MDL). In deployed fibers, the MDG or MDL generates a stochastic effect that turns the channel capacity into a random variable [11]. The stochastic capacity in space-division multiplexing (SDM) channels introduces a probability of outage if the receiver is set to operate on data rates higher than those instantaneously supported by the channel [12]. The effects of a stochastic capacity have been extensively addressed in the past in the scope of polarization-dependent loss (PDL) [13, 14, 15, 16]. While the loss in current optical devices for single-mode systems has small polarization dependency [17], current SDM devices have significant levels of MDG and MDL that cannot be neglected [18, 19, 20]. Communications systems with a stochastic signal-to-noise ratio (SNR) need to be designed to maintain the outage probability below accepted levels. To properly assess outages in the design of future SDM systems, statistical models for the capacity are necessary.
@@
The capacity loss statistics for a set of amplification schemes are discussed in [21]. Expressions for the variance of the per-mode capacity loss for four different amplification schemes are derived as a function of the channel MDL vector, discussed in [22]. A closed-form expression for the average capacity is presented in [23], but the capacity variance is not calculated. The impact of frequency diversity on the capacity distribution has been discussed in [24], but no analytical expression for the capacity distribution has been presented. The exact distribution for the capacity of a single-mode-fiber optical system with polarization-dependent gain (PDG) ($D=2$) has been presented in [25]. However, the statistics of SDM systems with higher mode counts have been addressed by simulation.
@@
In this paper, we provide analytical expressions for the variance of the capacity distribution [^1] An analytical solution for the total capacity mean is presented in [23, Eq. (10)].. The derivations are based on the assumption of a Gaussian distribution for the per-mode and total capacity [^2] The Gaussian assumption for the total capacity distribution has been discussed in [26], confirming its validity for increasing mode counts. probability density functions (PDFs). The joint distribution of per-mode capacities is approximated as a multivariate normal, and the total capacity is obtained as the linear combination of the per-mode capacities. The derivations rely on the modal gain distributions outlined in [23].
The remainder of this work is structured as follows. Section II reviews the channel capacity in strongly-coupled SDM systems subject to MDG and explains the most relevant statistical distributions that are investigated throughout the paper. Section III derives the statistical distribution of the per-mode capacity using two different methods. Section IV uses the previously defined per-mode capacity distribution to obtain the total capacity distribution. Section V investigates the effects of frequency diversity on the capacity variance and validates the derived analytical model. Section VI presents a step-by-step application of the derived model to estimate the outage capacity of a sample case study. Lastly, Section VII concludes the paper.
II Capacity Metrics In Optical SDM
Figure 1: Representation of a coupled SDM system with $N_{c}$ spectral channels and $D$ spatial channels. The spectral and spatial channels are first modulated. Subsequently, a spectral multiplexer generates a WDM signal, and a spatial multiplexer generates an SDM signal. The transmitted symbol sequences $s_{i}[n]$, for $i\in[1,D]$, are recovered independently for each spectral channel by a $D\times D$ MIMO equalization process, generating estimates $\widehat{s}_{i}[n]$ of the transmitted symbol sequence. The optical carrier modulator and the receiver front-end perform electrical-to-optical and optical-to-electrical conversion, respectively. The channel transfer matrix $\mathbf{H}$ and modal gains $\lambda_{\mathrm{dB}}$ are frequency dependent, but we omit the frequency dependent variable for simplicity, assuming a narrowband channel. The insets illustrate the evolution of the logarithmic scale modal gains PDF ($f_{\lambda_{\text{dB}}}(\lambda_{\text{dB}})$).
We consider the wavelength-division multiplexing (WDM) and SDM system presented in Fig. 1. The $N_{c}$-channel spectral multiplexer generates the WDM signal and the $D$-channel spectral multiplexer generates the SDM signal. In the mathematical derivations, we analyze only a narrowband channel, i.e., a channel with flat frequency response along the signal bandwidth. $D$ independent signals $s_{i}$, $i\in[1,D]$, are spatially multiplexed into the SDM fiber system for each frequency channel. The SDM signal is amplified multiple times by amplifiers with MDG that maintain a total output power evaluated across all spatial and spectral channels. Assuming a linear channel, transmission can be represented by a transfer matrix $\mathbf{H}$ followed by additive noise at the receiver [23]. The coupled signals are then separated at the receiver by a $D\times D$ multiple-input multiple-output (MIMO) digital signal processing (DSP) structure.
@@ II Capacity Metrics In Optical SDM
where SNR is the ratio of total signal power and total noise power, $\mathbf{H}$ is the channel transfer matrix, and $\mathbf{I}$ is the identity matrix of size $D$, where $D$ is the mode count. We assume a narrowband channel for most of the work. A discussion on the frequency dependence of the modal gains is provided in Section V. The total instantaneous capacity can also be obtained from the linear scale eigenvalues $\lambda$ of the $\mathbf{H}\mathbf{H}^{H}$ [27] operator
@@
where $\lambda_{i}$ is the $i^{th}$ Schmidt mode power gain [^3] The Schmidt modes are the optimal basis of propagation. For a discussion on Schmidt modes, see [28, 29].. We consider noise as spatially white [^4] The spatial whiteness assumption has stronger validity for higher mode counts and wider bandwidths [30]. The expressions presented in [21] also account for other amplification regimes., yielding equal noise power for all modes [23]. Assuming that SDM amplifiers maintain a constant total power after amplification, the sum of all $\lambda_{i}$ is set to $D$.
As shown in (2), the total capacity $C_{T}$ delivered by the system is the sum of the capacities of the $D$ individual Schmidt modes. The eigenvalues $\lambda_{i}$ in (2) can be interpreted as a relative multiplication factor that scales the SNR of the system for each mode. Without MDG, all $\lambda_{i}$s are the same. With MDG, they are different for each mode and vary over time. Let us now return to our primary question, which is to determine the distribution $f_{C_{T}}$ of the total capacity $C_{T}$. Having access to $f_{C_{T}}$ enables the system designer to avoid outages that could disrupt the system while minimizing resource overprovisioning. Clearly, $C_{T}$ depends on the distribution $f_{C_{i}}$ of the individual modes. The following derivations demonstrate that the distribution $f_{C_{i}}$ of each mode can be approximated as Gaussian. Consequently, as $C_{T}$ is the sum of the individual capacities $C_{i}$, it can also be assumed to follow a Gaussian distribution. Assuming both $C_{T}$ and $C_{i}$ as Gaussian, the distribution can be completely described by the mean and standard deviation. While the mean of $C_{T}$ has been previously derived in [24], its variance is derived for the first time in this paper. We first derive approximations for distribution $f_{\lambda_{\textrm{dB},i}}$ of $\lambda_{i}$, expressed in dB. Then, from $f_{\lambda_{\textrm{dB},i}}$ we obtain $f_{C_{i}}$ using properties of distribution transformation. Finally, assuming $f_{C_{i}}$ as Gaussian, we obtain a Gaussian $f_{C_{T}}$. This step uses conventional properties of sums of correlated Gaussian random variables.
Fig. 2 illustrates the most important distributions addressed in this paper. Fig. 2a presents the distribution of the power gains $\lambda_{i}$ in logarithmic scale. The distributions for the individual $\lambda_{i}$s, shown by markers, are obtained by Monte-Carlo simulations [^5] The simulation results rely on the multisection model discussed in [23]. The per-section MDG is controlled such that a target $\sigma_{\mathrm{mdg}}$ is achieved. From the channel matrix, we obtain the instantaneous modal gains and capacity values. The statistical distribution and values are obtained over 100 Monte-Carlo realizations. The results are obtained by simulating 100 sections with 50 km each. The mode count, SNR, and $\sigma_{\mathrm{mdg}}$ values of each simulation are specified on the respective figures.. The solid black curve represents the analytical distribution of the ensemble of power gains $f_{\lambda_{\textrm{dB}}}$ as presented in [23], normalized by the number of modes, and is included here for reference. In a later section, we derive an analytical approximation for this distribution for the first time. Fig. 2b illustrates the distribution of the per-mode capacity $C_{i}$ and the total capacity $C_{T}$. Again, these distributions are obtained through simulations, and analytical expressions are derived in later sections. It is interesting to see how the different modes contribute to different capacities. The figures also suggest that the distribution of $C_{i}$ and $C_{T}$ can be approximated as Gaussian.
Figure 2: (a) Logarithmic gain $\lambda_{\mathrm{dB}}$ distribution of each mode and ensemble, considering the sum of the linear gains equals to $D$. The ensemble distribution is normalized by $D$ in order to facilitate visual analysis. The per-mode distributions are obtained via simulation for ${D=6}$ and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$. (b) Instantaneous capacity distribution of each mode and total. The distributions are obtained via simulation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$. Dotted lines connect the simulated histogram values for visualization purposes.
III Per-Mode Capacity Distribution $C_{i}$
@@
We evaluate the statistics of the total capacity $C_{T}$ based on the per-mode individual $\lambda_{i}$ and $C_{i}$ distributions. Approximations on $C_{i}$ distributions are obtained from two methods derived from the Gaussian unitary ensemble (GUE) spectral and Wigner semicircular distributions for $\lambda_{i}$.
@@ III-A Approximations on the $C_{i}$ distribution based on the GUE spectral distribution for $\lambda_{i}$
The PDF $f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})$ of the $\lambda_{\mathrm{dB}}$ eigenvalues for a $D^{th}$ order GUE is given by [23, Table 1], [31, Eq. (S.15)]
@@
where $\alpha_{\lambda_{\mathrm{dB}},D}$ is a normalization factor such that the PDF integral equals 1, and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are polynomial coefficients. $\alpha_{\lambda_{\mathrm{dB}},D}$ and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are constant for a given number of modes. For $D\leq 8$ the values of $\alpha_{\lambda_{\mathrm{dB}},D}$ and $\beta_{\lambda_{\mathrm{dB}},D,k}$ are provided in [23, Table 1], and the derivation of the coefficients is presented in appendix A. Values [^6] The notation $\sigma_{\mathrm{mdg}}$ was preferred to $\sigma_{\lambda_{\mathrm{dB}}}$, as the former is predominant in literature. $\sigma_{\mathrm{mdg}}$ and $\mu_{\lambda_{\mathrm{dB}}}$ are the standard deviation and the mean of $f_{\lambda_{\mathrm{dB}}}(\lambda_{\mathrm{dB}})$, respectively. Value $\mu_{\lambda_{\mathrm{dB}}}$ is given such that the linear scale eigenvalue distribution $f_{\lambda}(\lambda)$ has unitary mean, and is given by (see appendix B)
@@
Figure 3: Analytical approximation of per-mode capacity mean (a) and standard deviation (b) for ${\mathrm{SNR}=10~{}\mathrm{dB}}$ and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$. For ${D\leq 8}$ the GUE distribution method is considered. Higher mode counts use the Wigner semicircular CDF method.
Figure 4: Comparison between the analytically obtained curves, assuming a Gaussian distribution, and simulation values for the per-mode capacity PDF s for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$.
@@
An evaluation of $\mu_{C_{i}}$ and $\sigma_{C_{i}}$ for $2\leq D\leq 8$ is presented in Fig. 3. The accuracy of the Gaussian approximation for the individual capacity distributions $f_{C_{i}}$, for a system with $D=6$, $\mathrm{SNR}=10~{}\mathrm{dB}$ and $\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}$, is shown in Fig. 4. The results reveal, again, a suitable agreement between simulation and the derived analytical curves.
@@
The limiting distribution of $\lambda_{\mathrm{dB}}$ as $D\rightarrow\infty$ is given by the Wigner semicircular distribution as [32, I][23, Table 1]
@@
An assessment of the CDF-derived approximation can be found in Fig. 5, for a system with $D=6$, $\mathrm{SNR}=10~{}\mathrm{dB}$ and $\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}$, showing excellent agreement.
@@
IV Total Capacity Distribution $C_{T}$
@@
The average $\mu_{C_{T}}$ presented in (18) holds true given our assumption that the per-mode capacity is Gaussian distributed. Alternatively, an exact solution for $\mu_{C_{T}}$, derived directly from the modal gains ensemble distribution, has already been proposed in [23, Eq. (10)][25, Eq. (12)]
@@
Figure 6: Correlation for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$. Dotted lines connect the analytical correlation values for visualization purposes.
@@
$\rho_{i,j}=e^{-|i-j|}+\left(e^{-|i-j|}-1\right)\left(\gamma_{0}+\gamma_{1}\cdot\sigma_{\mathrm{mdg}}^{2.75}\right),$ (21)
@@
$\left(\gamma_{0},\gamma_{1}\right)=\mathrm{argmin}_{\gamma_{0},\gamma_{1}}\left(\mathbb{E}\left\{\left[\mathrm{ln}\left(\widehat{\sigma}^{2}_{C_{T}}\right)-\mathrm{ln}\left(\sigma^{2}_{C_{T}}\right)\right]^{2}\right\}\right),$ (22)
where $\widehat{\sigma}^{2}_{C_{T}}$ is the total capacity variance obtained via simulation.
The methodology to obtain the coefficients is performed iteratively until a tolerance threshold is reached. The algorithm implemented in this work firstly updates $\gamma_{0}$ until the error for the lowest considered $\sigma_{\mathrm{mdg}}$ is smaller than a defined tolerance. The value of $\gamma_{1}$ is updated considering the MSLE for the entire evaluated $\sigma_{\mathrm{mdg}}$ range, with small corrections to $\gamma_{0}$ if the resulting approximation function is not monotonic increasing. We select the logarithmic error because the capacity variance scales exponentially with $\sigma_{\mathrm{mdg}}$ when measured in decibels. The coefficients for a practical range of mode counts and SNR are given in Table I.
An evaluation of the analytical method is presented in Fig. 7. The results indicate an accurate variance approximation for SNR values lower than $20~{}\mathrm{dB}$ for all evaluated mode counts. For $\textrm{SNR}=20~{}\mathrm{dB}$, there is a higher divergence, particularly with the Wigner semicircular CDF method. The source of estimation discrepancy likely stems from the per-mode individual distributions estimated from the capacity CDF, as well as approximations made in the correlation function. Fig. 8 evaluates the accuracy of the method based on Wigner semicircular CDF for different values of $\sigma_{\mathrm{mdg}}$ and mode counts. The results evidence accurate estimates for mode counts higher than 20. Fig. 9 compares the simulated total capacity distribution with the analytical distributions produced with the Wigner semicircular CDF method. The presented Gaussian approximation exhibits a good agreement with the simulated curve.
Figure 9: Total Gaussian PDF for ${D=20}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$.